# Restricted Path Characteristic Function Determines the Law of Stochastic Processes

### v2

@@ Abstract.
A central question in rough path theory is characterising the law of stochastic processes. It is established in [I. Chevyrev &\&& T. Lyons, Characteristic functions of measures on geometric rough paths, Ann. Probab. 44 (2016), 4049–4082] that the path characteristic function (PCF), i.e., the expectation of the unitary development of the path, uniquely determines the law of the unparametrised path. We show that PCF restricted to certain subspaces of sparse matrices is sufficient to achieve this goal. The key to our arguments is an explicit algorithm — as opposed to the nonconstructive approach in [I. Chevyrev &\&& T. Lyons, op. cit.] — for determining a generic element $X$ of the tensor algebra ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}=\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$ from its moment generating function. Our only assumption is that $X$ has a nonzero radius of convergence, which relaxes the condition of having an infinite radius of convergence in the literature.
As applications of the above theoretical findings, we propose the restricted path characteristic function distance (RPCFD), a novel distance function for probability measures on the path space that offers enormous advantages for dimension reduction. Its effectiveness is validated via hypothesis testing on fractional Brownian motions, thus demonstrating the potential of RPCFD in generative modeling for synthetic time series generation.
@@ 1. Introduction
Charaterisation of the law of stochastic processes is of enormous significance in stochastic analysis, statistics, and machine learning [16]. The study of faithful and parsimonious representations of laws on the path space, in particular, has a direct impact on algorithm design and efficiency. It has found successful applications in numerical methods of SDE simulations, hypothesis testing on stochastic processes, and generative models for synthetic time series, etc. [25, 24, 23, 26].
@@
The signature of paths, a central concept in rough path theory, can be regarded as non-commutative monomials on the path space. It uniquely determines the underlying path of finite $p$-variations up to a negligible equivalent class; $p\geq 1$. Therefore, characterising the law of stochastic processes is equivalent to characterising the measures on the space of path signatures. It is shown in [7] that the path characteristic function (PCF), i.e., the expectation of the unitary development of the path, determines the law of the signature.
In this paper, we prove that a sparse version of PCF determines the law of the signature under very mild regularity conditions, thus significantly enhancing the results in [7]. Our approach is the natural analogue in the (infinite-dimensional) tensor algebra setting of the classical moment problem, i.e., determination of the law of an $\mathbb{R}$-valued random variable from its moment generating function.
To be more precise, consider
$X=S\big{(}\gamma_{[0,T]}\big{)},$
where $\gamma$ is a stochastic process on $\mathbb{R}^{d}$ and $S$ denotes the signature map. Assume that $\gamma$ is of bounded $p$-variations with $21\leq p<$ 2, so that $X$ takes values of group-like elements in $\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$. The key results in Chevyrev–Lyons [7, Theorem 4.8; Corollary 4.12] ascertain that the characteristic function (see (12) for definition) over $\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$-valued linear mappings determines the law of $X$, and hence essentially [^1] That is, modulo tree-like equivalence; cf. [10] that of $\gamma_{[0,T]}$. Our paper improves [7] by showing that the same conclusion remains valid when $\bigoplus_{k=0}^{\infty}\mathfrak{su}(k)$ above is replaced by the subalgebra $\bigoplus_{k=0}^{\infty}\mathfrak{so}(k;\mathbb{R})$, or even by the subalgebra of very sparse tridiagonal matrices $\bigoplus_{k=0}^{\infty}\mathfrak{t}(k,\mathbb{R})$ in (1) below. In machine learning applications, this provides enormous benefits to dimension reduction in Cartan development-based distance functions for measures on the path space. See §§5 and 6.
In a much broader context, we study the “generalised moment problem”:
Determine a tensor $X\in\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$ with a nonzero radius of convergence (see §1.4) from its moment generating function.
The only requirement is that the radius of convergence is nonzero, which relaxes the condition of infinite radius of convergence in [7]. We establish an explicit, constructive algorithm for the generalised moment problem in the tensor algebra. The Main Theorem 1.8 states that given the unitary development of an unknown tensor with nonzero radius of convergence, by selecting a few sparse matrices according to certain combinatorial rules, taking the Cartan development of their linear combination, differentiating with respect to the coefficients (“parameters”) $\theta$ in the linear combination, and finally evaluating at $\theta=0$, we are able to recover the tensor.
Our algorithm justifies that the unitary development constitutes a good feature representation of paths in machine learning applications, which will be substantiated in §6. Most notably, the matrices involved in the algorithm are rather sparse: they take values in
$\mathfrak{t}(k,\mathbb{R}):=\bigg{\{}M\in\mathbb{R}^{k\times k}:\,\text{$M+M^{\top}=0$, and the $(i,j)$-entry of $M$}\\ \text{is possibly non-zero only when $|i-j|=1$}\bigg{\}},$ (1)
hence rendering considerable advantages in dimension reduction.
As a corollary, if a tensor $X$ with positive or infinite radius of convergence is nonzero at the $n^{\text{th}}$ level, then for each $k\geq n+1$ one can find a linear map $M:\mathbb{R}^{d}\to\mathfrak{t}(k,\mathbb{R})$ such that the unitary development of $X$ evaluated at $M$ is nonzero. In contrast, [7, Theorem 4.8] establishes that if a tensor $X$ with infinite radius of convergence is nonzero at the $n^{\text{th}}$ level, one can find $M:\mathbb{R}^{d}\to\mathfrak{sp}(m)$ for some $m\geq\max\left\{2,\frac{n}{3}\right\}$ such that the unitary development of $X$ evaluated at $M$ is nonzero. Here $\mathfrak{sp}(m)$, the Lie algebra whose complexification is the symplectic matrix algebra of size $2m\times 2m$, contains $\mathfrak{t}(2m,\mathbb{R})$ as a subspace. The latter result is proved in [7] via polynomial identities in numerical linear algebra works [8, 9] in a non-constructive manner. It only ascertains the existence of such $M$, but gives no clue at all on explicit constructions of $M$.
@@
To describe our results in full rigour, let us first introduce the notions of signature, development, generating function, and radius of convergence in the subsections below.
@@ 1.1. Path signature
Consider a path $\gamma:[0,T]\to\mathbb{R}^{d}$ of bounded $p$-variation; $p\in[1,2)$. The signature of $\gamma$, first introduced by K.-T. Chen in 1950’s ([3, 4]), is defined as a sequence of iterated integrals:
$S(\gamma)_{[0,T]}:=\left(1,\gamma^{(1)}_{[0,T]},\ldots,\gamma^{(n)}_{[0,T]},\ldots\right),$
$\gamma^{(n)}_{[0,T]}:=\int_{0<t_{1}<t_{2}<\ldots<t_{n}<T}{\rm d}\gamma_{t_{1}}\otimes\cdots\otimes{\rm d}\gamma_{t_{n}}\,\in\left(\mathbb{R}^{d}\right)^{\otimes n}$ (2)
that takes values in the tensor algebra:
${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}:=\bigoplus_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n},$ (3)
with the convention $\left(\mathbb{R}^{d}\right)^{\otimes 0}\equiv\mathbb{R}$. Here the iterated integral in (2) is understood in Young’s sense [17]. It is easy to see that the signature is invariant under time-reparametrisation; hence, we shall drop the subscript [0,T] from $S(\gamma)$ from now on.
The mapping $\gamma\mapsto S(\gamma)$ defines a group homomorphism from the path group (with concatenation $\ast$ and path-inverse $\overleftarrow{\bullet}$ as group operation and inverse) into ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, such that
@@
The signature contains all essential information of a path. More precisely, Chen showed in [5] that two irreducible piecewise $C^{1}$-paths cannot have the same signature unless they differ only by a translation or reparametrisation. This result was extended to BV-paths by Hambly–Lyons in [10], proving that the signature uniquely determines BV-paths modulo tree-like equivalence.
@@ 1.2. Path development
${\mathscr{D}}_{\gamma}:{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)\longrightarrow G$
@@
where $I_{k}$ is the identity of $G$. This is well defined when $\gamma$ is a path of bounded $p$-variation; $21\leq p<$ 2. The symbol $\bullet$ denotes matrix multiplication, and ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)$ is the space of linear transforms from $\mathbb{R}^{d}$ to ${\mathfrak{g}}$. The formula (5) is closely related to the signature of $\gamma$ defined in (2).
@@
For $G=U(k)=$ the unitary group and ${\mathfrak{g}}={\mathfrak{su}}(k)=$ the unitary algebra (i.e., the space of skew-Hermitian matrices $M$ with $M+M^{*}=0$), we call ${\mathscr{D}}_{\gamma}$ the unitary development of $\gamma$. In machine learning papers (e.g., [15, 14]), it is also known as the “unitary feature”.
@@ 1.3. Generating function of a tensor
Let $x\equiv\{x_{n}\}_{n=0}^{\infty}$ be a real sequence. Its generating function is given by the (formal) complex series:
$\phi_{x}:\mathbb{C}\longrightarrow\mathbb{C},\qquad\phi_{x}(\lambda)=\sum_{n=0}^{\infty}x_{n}\lambda^{n}.$ (6)
@@
For each given $k=1,2,3,\ldots$ and tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, the generating function of $X$ of order $k$ is the mapping
$\Phi_{X}^{(k)}:{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\,\longrightarrow\,{\mathfrak{gl}(k,\mathbb{C})},$
$\Phi_{X}^{(k)}(M):={\widetilde{M}}(X).$ (8)
@@
${\widetilde{M}}$ is the algebra homomorphism determined by ${\widetilde{M}}(1)=I_{k}$ and that
${\widetilde{M}}\left(e_{i_{1}}\otimes\cdots\otimes e_{i_{\ell}}\right):=M\left(e_{i_{1}}\right)\bullet\cdots\bullet M\left(e_{i_{\ell}}\right)\text{ for any }\ell\in\mathbb{N}\text{ and }i_{1},\ldots,i_{\ell}\in\{1,\ldots,d\}.$ (9)
We also denote the generating function of $X$ as
$\Phi_{X}:\bigoplus_{k=0}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\,\longrightarrow\,\bigoplus_{k=0}^{\infty}{\mathfrak{gl}(k,\mathbb{C})},$
$\Phi_{X}\Big{|}_{{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)}=\Phi_{X}^{(k)},$
which respects the natural grading in $k\in\mathbb{N}$.
We define ${\widetilde{M}}$ by (1.1) since
$\{1\}\cup\big{\{}e_{i_{1}}\otimes\cdots\otimes e_{i_{\ell}}:\ell\in\mathbb{N}\text{ and }i_{1},\ldots,i_{\ell}\in\{1,\ldots,d\}\big{\}}$
forms a Hamel basis for ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$. Notice too that the defining equation (1.1) for the generating function is formal — consider the projection (a.k.a. truncation) of a tensor onto the $n^{\text{th}}$ level:
$\pi_{n}:{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}\to\left(\mathbb{R}^{d}\right)^{\otimes n}.$ (10)
In view of (1.1), the action of ${\widetilde{M}}$ on $\pi_{n}(X)$ involves a product of $n$ matrices. To make (1.1) rigorous, we require that each $n$-fold product in (1.1) converges in ${\mathfrak{gl}(k,\mathbb{C})}$ as the tensor degree $n\to\infty$. To fix the idea, in this paper we always adopt the Hilbert–Schmidt norm on ${\mathfrak{gl}(k,\mathbb{C})}$.
Definition 1.1 of the generating function can be extended to any Lie algebra $\mathfrak{g}$ and normed unital algebra $\mathfrak{a}$. One defines $\Phi_{X}:\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{g}\right)\to\mathfrak{a}$ again via $\Phi_{X}(M):={\widetilde{M}}(X)$, where ${\widetilde{M}}$ is given by (1.1) with $\bullet$ being the multiplication in $\mathfrak{a}$. If $X\in\mathcal{G}\left(\mathbb{R}^{d}\right)$, then $\Phi_{X}$ takes values in $G$, the Lie group of $\mathfrak{g}$. On the other hand, if $X$ is in ${\rm Lie}\left[\mathcal{G}\left(\mathbb{R}^{d}\right)\right]$, the Lie algebra of $\mathcal{G}\left(\mathbb{R}^{d}\right)$, then $\Phi(X)\in\mathfrak{g}$ [^2] The Lie algebra of $\mathcal{G}\left(\mathbb{R}^{d}\right)$ is denoted as $\mathcal{L}((\mathbb{R}^{d}))$ in [17]. We avoid using this notation and reserve $\mathcal{L}$ for the space of linear maps. Here and hereafter,
@@
whose Lie algebra ${\rm Lie}\left[\mathcal{G}\left(\mathbb{R}^{d}\right)\right]$ is spanned by $\left[e_{i_{1}},\left[e_{i_{2}},\cdots,\left[e_{i_{\ell-1}},e_{i_{\ell}}\right]\right]\right]$ where $\{e_{1},\ldots,e_{d}\}$ is a basis for $\mathbb{R}^{d}$. See [17, §2.2.4].
@@
${\mathcal{T}}(\mathbb{R})\cong\bigoplus_{n=0}^{\infty}\mathbb{R}=\text{the space of real sequences}.$
@@
Observe the structural resemblance of the integrand of the development in (5) with the canonical extension in (1.1) of the generating function. Indeed, we have the following result.
@@
The development and signature of a BV-path $\gamma:[0,T]\to\mathbb{R}^{d}$ are related by the following: for any $k=1,2,3,\ldots$ and $M\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$, it holds that
${\mathscr{D}}_{\gamma}(M)={\widetilde{M}}\big{(}S(\gamma)\big{)}.$
Proof.
See [14, Appendix A]; ${\mathfrak{su}}(k)$ can be replaced by any finite-dimensional Lie algebra. ∎
@@
In light of Proposition 1.3, if $X=S(\gamma)$ for some random BV-path $\gamma$, then $\phi_{X}(M)$ equals the expectation of the development ${\mathscr{D}}_{\gamma}(M)$. Here $G$ can be taken as $\mathcal{G}\left(\mathbb{R}^{d}\right)$, the group-like elements in the tensor algebra ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$; cf. [7, §3].
1.4. The generalised moment problem; radius of convergence
One central question concerning the (unitary) development is whether the analogue of (7) holds in the case of tensors.
Question 1.4.
Given the mapping $\Phi:\bigoplus_{k=0}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\to\bigoplus_{k=0}^{\infty}{\mathfrak{gl}(k,\mathbb{C})}$ graded by $k\in\mathbb{N}$, which is known to be the generating function of some $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, can one solve for $X$?
We shall refer to it as the generalised moment problem [^3] Indeed, in Chevyrev–Lyons [7, §6.1] “the moment problem” refers to our generalised moment problem specialised to the case of expected signatures. for $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$.

### v3

@@ Abstract.
A central question in rough path theory is characterising the law of stochastic processes on path spaces. It is established in [I. Chevyrev &\& T. Lyons, Characteristic functions of measures on geometric rough paths, Ann. Probab. 44 (2016), 4049–4082] that the characteristic function of a probability measure on group-like elements, which is a subspace of the extended tensor algebra ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}=\prod_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$, uniquely determines the measure. In this work, we show that the characteristic function restricted to special orthogonal Lie algebra $\mathfrak{so}$}(n) is sufficient to achieve this goal. The key to our arguments is an explicit algorithm — as opposed to the non-constructive approach in [I. Chevyrev &\& T. Lyons, op. cit.] — for determining a generic element $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ from its generating function when restricting its domain to a sparse subspace of real tridiagonal skew-symmetric matrices. Our only assumption is that $X$ has a non-zero ROC, which relaxes the condition of having an infinite ROC in the literature.
As an application, we propose the restricted path characteristic function distance (RPCFD), a novel distance function for probability measures on the path space that serves as the sparse counterpart of path characteristic function distance. It has enormous advantages in dimension reduction and potential in generative modeling for synthetic time series generation, validated in this paper via hypothesis testing on fractional Brownian motions.
@@ 1. Introduction
Characterising the law of stochastic processes is of enormous significance in stochastic analysis, statistics, and machine learning [17]. The study of faithful and parsimonious representations of laws on the path space, in particular, has a direct impact on algorithm design and efficiency. It has found successful applications in numerical methods of SDE simulations, hypothesis testing on stochastic processes, and generative models for synthetic time series, etc. [27, 26, 25, 28].
@@
The signature of paths, a central concept in rough path theory, can be regarded as non-commutative monomials on the path space. It uniquely determines the underlying path of finite $p$-variations up to a negligible equivalence class where $p\geq 1$ (i.e., the tree-like equivalence [10, 11]). Therefore, characterising the law of stochastic processes is equivalent to characterising probability measures on the space of path signatures.
In this paper, we consider measures on some subspace $\mathscr{V}$ (to be specified later) larger than the space of path signatures in the extended tensor algebra ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}:=\prod_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n}$.Let $\mu$ be a Radon measure on $\mathscr{V}$. We define its characteristic function by
$\Phi_{\mu}(M):=\mathbb{E}_{X\sim\mu}\left[\widetilde{M}(X)\right].$ (1)
Here, $M$ ranges over linear mappings from $\mathbb{R}^{d}$ to $\bigoplus_{k=1}^{\infty}\mathfrak{su}(k)$, i.e., the space of skew-Hermitian matrices, and $\widetilde{M}$ denotes the canonical extension of $M$. See (14) for its precise definition. Such a generalisation of the characteristic function appeared in [7], and it coincides with the path characteristic function PCF, namely the expectation of the unitary development of the stochastic process, when the measure $\mu$ is restricted to the space of path signatures [16].
The key result of the seminal paper [7] by Chevyrev–Lyons ascertains that PCF determines the law of the signature of the underlying stochastic process, and PCF is determined by the expected signature provided that the latter has an infinite radius of convergence (ROC). The proof in [7] is non-constructive and leaves open two important questions: (1), whether the entire PCF, rather than some of its simpler variants, is necessary in order to determine the law of signature, and (2), how the PCF or its variants determine the expected signature, if the expected signature exists and satisfies certain regularity conditions. Our work offers a novel and constructive approach to address these questions. We show that PCF, when restricted to $\bigoplus_{k=1}^{\infty}\mathfrak{so}(k)$ (i.e., the space of real skew-symmetric matrices), is sufficient to determine the law of signature. For the case where the expected signature has the infinite radius of convergence, our approach provides an explicit way to recover the expected signature, which can not be achieved by the approach of [7].
To achieve these results, a key problem is to show that the family of matrix-valued functions
$\mathcal{A}:=\left\{\widetilde{M},\text{ canonical extension of }M\in\bigoplus_{k=1}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{su}(k)\right)\right\}$ (2)
separates points over $\mathscr{V}\subset{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, in view of the Stone–Weierstrass theorem. Chevyrev–Lyons established in [7, Theorem 4.8] that $\mathcal{A}$ separates points over the subspace of tensors with infinite radii of convergence (ROC) by exploiting polynomial identities in the numerical linear algebra literature [8, 9]. Our Main Theorem 1.9 strengthens this result in three aspects:
  1. (1)
    We show that $\mathcal{A}$ indeed separates points over the subspace of tensors with non-zero ROC.
  2. (2)
    We show that a (much) smaller subset $\mathcal{A}_{\mathfrak{t}}\subset\mathcal{A}$ is sufficient to achieve this goal.
  3. (3)
    We give an explicit algorithm for recovering $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$ from $\left\{\widetilde{M}(X):\,{\widetilde{M}}\in\mathcal{A}_{\mathfrak{t}}\right\}$, provided that $X$ has a non-zero ROC.
Here and hereafter, we set
$\mathcal{A}_{\mathfrak{t}}:=\left\{\widetilde{M},\text{ canonical extension of }M\in\bigoplus_{k=1}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};\mathfrak{t}(k,\mathbb{R})\right)\right\},$
where $\mathfrak{t}(k,\mathbb{R})$ is the space of “tridiagonal” skew-symmetric real matrices:
$\mathfrak{t}(k,\mathbb{R}):=\bigg{\{}M\in\mathbb{R}^{k\times k}:\,\text{$M+M^{\top}=0$; the $(i,j)$-entry of $M=0$ whenever $|i-j|\neq 1$}\bigg{\}}.$ (3)
@@
Let us first introduce the notions of signature, development, generating function, and radius of convergence in the subsections below.
@@ 1.1. Path signature
Consider a path $\gamma:[0,T]\to\mathbb{R}^{d}$ of bounded $p$-variations where $1\leq p<2$. The signature of $\gamma$, first introduced by K.-T. Chen in 1950’s ([3, 4]), is defined by a sequence of iterated integrals:
[…]
$\gamma^{(n)}_{[0,T]}$ $:=\int_{0<t_{1}<t_{2}<\ldots<t_{n}<T}{\rm d}\gamma_{t_{1}}\otimes\cdots\otimes{\rm d}\gamma_{t_{n}}\,\in\left(\mathbb{R}^{d}\right)^{\otimes n},$ (5)
that takes values in the extended tensor algebra:
${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}:=\prod_{n=0}^{\infty}\left(\mathbb{R}^{d}\right)^{\otimes n};$ (6)
set $\left(\mathbb{R}^{d}\right)^{\otimes 0}\equiv\mathbb{R}$. The iterated integral in (5) is understood in Young’s sense [18]. The signature is time-reparametrisation-invariant; hence, we drop the subscript [0,T] in $S(\gamma)$ from now on.
[…]
@@
The signature contains all essential information of a path. More precisely, Chen showed in [5] that two irreducible piecewise $C^{1}$-paths cannot have the same signature unless they differ only by a translation or reparametrisation. This result was extended to BV-paths by Hambly–Lyons in [10], proving that the signature uniquely determines BV-paths modulo tree-like equivalence. And we refer to [11] for the tree-like equivalence of finite $p$-variation paths where $1<p<2$.
@@ 1.2. Path development
${\mathscr{D}}_{\gamma}:{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)\longrightarrow G$ (8)
@@
where $I_{k}$ is the identity of $G$. This is well defined when $\gamma$ is a path of bounded $p$-variation for $1\leq p<2$. The symbol $\bullet$ denotes matrix multiplication, and ${\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)$ is the space of linear transforms from $\mathbb{R}^{d}$ to $\mathfrak{g}$}. The formula (9) is closely related to the signature of $\gamma$ defined in (5).
@@
For $G=U(k)=$ the unitary group and ${\mathfrak{g}}={\mathfrak{su}}(k)=$ the unitary algebra (i.e., the space of skew-Hermitian matrices $M$ with $M+M^{*}=0$), we call ${\mathscr{D}}_{\gamma}$ the unitary development of $\gamma$. In machine learning papers (e.g., [16, 15]), it is also known as the unitary feature.
Remark 1.1.
Let $1\leq p<2$. The development of a finite $p$-variation path $\gamma:[0,T]\rightarrow\mathbb{R}^{d}$ under $M\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{g}}\right)$ can be equivalently defined by the unique solution of the following differential equation
${\rm d}Y_{t}=Y_{t}\bullet M({\rm d}\gamma_{t}),\quad Y_{0}=I_{k},$ (10)
where the integral is understood in Young’s sense. Notice that (9) coincides with the Picard iteration of (10).
@@ 1.3. Generating function of a tensor
Let $x\equiv\{x_{n}\}_{n=0}^{\infty}$ be a real sequence. Its generating function is defined as the (formal) complex series:
$\phi_{x}:\mathbb{C}\longrightarrow\mathbb{C},\qquad\phi_{x}(\lambda):=\sum_{n=0}^{\infty}x_{n}\lambda^{n}.$ (11)
@@
For each given $k=1,2,3,\ldots$ and tensor $X\in{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$, the generating function of $X$ of order $k$ is defined by the mapping:
$\Phi_{X}^{(k)}:{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\,\longrightarrow\,{\mathfrak{gl}(k,\mathbb{C})},\qquad\Phi_{X}^{(k)}(M):={\widetilde{M}}(X).$ (13)
@@
[…]
The generating function of $X$ is defined by:
$\Phi_{X}:\bigoplus_{k=1}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\,\longrightarrow\,\bigoplus_{k=1}^{\infty}{\mathfrak{gl}(k,\mathbb{C})},$ (15)
$\Phi_{X}\Big{|}_{{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)}:=\Phi_{X}^{(k)},$ (16)
which respects the natural grading in $k\geq 1$.
Notice that the defining equation (13) for the generating function is formal — consider the projection of a tensor onto the $n^{\text{th}}$ level:
$\pi_{n}:{{\mathcal{T}}\left(\mathbb{R}^{d}\right)}\longrightarrow\left(\mathbb{R}^{d}\right)^{\otimes n}.$ (17)
In view of (14), the action of $\widetilde{M}$} on $\pi_{n}(X)$ involves a product of $n$ matrices. To make (13) rigorous, we require that each $n$-fold product in (14) converges in $\mathfrak{gl}(k,\mathbb{C})$} as the tensor degree $n\to\infty$. To fix the idea, in this paper we always adopt the Hilbert–Schmidt norm on $\mathfrak{gl}(k,\mathbb{C})$}.
Definition 1.2 of the generating function can be extended to any Lie algebra $\mathfrak{g}$ and normed unital algebra $\mathfrak{a}$. One defines $\Phi_{X}:\mathcal{L}\left(\mathbb{R}^{d};\mathfrak{g}\right)\to\mathfrak{a}$ again via $\Phi_{X}(M):={\widetilde{M}}(X)$, where $\widetilde{M}$} is given by (14) with $\bullet$ being the multiplication in $\mathfrak{a}$. If $X\in\mathcal{G}\left(\mathbb{R}^{d}\right)$, then $\Phi_{X}$ takes values in $G$, the Lie group of $\mathfrak{g}$. On the other hand, if $X$ is in ${\rm Lie}\left[\mathcal{G}\left(\mathbb{R}^{d}\right)\right]$, the Lie algebra of $\mathcal{G}\left(\mathbb{R}^{d}\right)$, then $\Phi(X)\in\mathfrak{g}$. [^1] The Lie algebra of $\mathcal{G}\left(\mathbb{R}^{d}\right)$ is denoted as $\mathcal{L}((\mathbb{R}^{d}))$ in [18]. We avoid using this notation and reserve $\mathcal{L}$ for the space of linear maps. Here and hereafter,
@@
and ${\rm Lie}\left[\mathcal{G}\left(\mathbb{R}^{d}\right)\right]$ is spanned by $\left[e_{i_{1}},\left[e_{i_{2}},\cdots,\left[e_{i_{\ell-1}},e_{i_{\ell}}\right]\right]\right]$ where $\{e_{1},\ldots,e_{d}\}$ is a basis for $\mathbb{R}^{d}$. See [18, §2.2.4].
@@
${\mathcal{T}}(\mathbb{R})\cong\prod_{n=0}^{\infty}\mathbb{R}=\text{the space of real sequences}.$
@@
Observe the structural resemblance of the integrand of the development in (9) with the canonical extension in (14) of the generating function. Indeed, we have the following result, in which ${\mathfrak{su}}(k)$ can be replaced by any finite-dimensional Lie algebra.
@@
Let $1\leq p<2$. The development and signature of a finite $p$-variation path $\gamma:[0,T]\to\mathbb{R}^{d}$ satisfies
${\mathscr{D}}_{\gamma}(M)={\widetilde{M}}\big{(}S(\gamma)\big{)}$
for any $k=1,2,3,\ldots$ and $M\in{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)$.
@@
In light of Proposition 1.4, if $X=S(\gamma)$ for some random finite $p$-variation path $\gamma$ where $1\leq p<2$, then $\phi_{X}(M)$ equals the expectation of the development ${\mathscr{D}}_{\gamma}(M)$. Here $G$ can be taken as $\mathcal{G}\left(\mathbb{R}^{d}\right):=\left\{\text{group-like elements in ${{\mathcal{T}}\left(\mathbb{R}^{d}\right)}$}\right\}.$ See [7, §3].
1.4. Radius of convergence of a tensor
One central question concerning the generating function of a tensor is whether the analogue of (12) holds.
Question 1.5.
Given $\Phi_{X}:\bigoplus_{k=1}^{\infty}{\mathcal{L}}\left(\mathbb{R}^{d};{\mathfrak{su}}(k)\right)\to\bigoplus_{k=1}^{\infty}{\mathfrak{gl}(k,\mathbb{C})}$ the generating function of $X$, graded by $k\geq 1$. Can we solve for $X$?